convert atm pressure from barometer reading in inches Hg to mm Hg and atm. given. 1atm= 760mmHg 1inche=2.54cm 1mm=0.1cm
1
Expert's answer
2017-03-21T02:52:06-0400
1inche=2.54cm = 25.4 mm 1 atm = 760mmHg / 25.4 mm = 29.921373 inHg
1 atm is the pressure equal to the weight of a mercury column 760 mm high at a temperature of 0.0 ° C, at a latitude of 45 ° and at sea level.
1 mm Hg = 101325/760 Pa. Formula for translation units is as follows:
Pp = PHg * 101325/760,
Where PHg - pressure, expressed in millimeters of mercury, Pp - pressure expressed in Pascal. Pp = 760mmHg * 101325/760mmHg = 101325 Pa or 1 atm